### v1

# Only a Little to the Left: A Theory-grounded Measure of Political Bias in Large Language Models

Abstract
Prompt-based language models like GPT4 and LLaMa have been used for a wide variety of use cases such as simulating agents, searching for information, or for content analysis. For all of these applications and others, political biases in these models can affect their performance. Several researchers have attempted to study political bias in language models using evaluation suites based on surveys, such as the Political Compass Test (PCT), often finding a particular leaning favored by these models. However, there is some variation in the exact prompting techniques, leading to diverging findings and most research relies on constrained-answer settings to extract model responses. Moreover, the Political Compass Test is not a scientifically valid survey instrument. In this work, we contribute a political bias measured informed by political science theory, building on survey design principles to test a wide variety of input prompts, while taking into account prompt sensitivity. We then prompt 11 different open and commercial models, differentiating between instruction-tuned and non-instruction-tuned models, and automatically classify their political stances from 88,110 responses. Leveraging this dataset, we compute political bias profiles across different prompt variations and find that while PCT exaggerates bias in certain models like GPT3.5, measures of political bias are often unstable, but generally more left-leaning for instruction-tuned models.
Only a Little to the Left: A Theory-grounded Measure of Political Bias in Large Language Models
Mats Faulborn1, Indira Sen2, Max Pellert3, Andreas Spitz1, and David Garcia1 1University of Konstanz 2University of Mannheim 3Barcelona Super Computing Center Correspondence: David.Garcia@uni-konstanz.de
1 Introduction
In the past years, Large Language Models (LLMs) have emerged as a transformative technology with applications in a large variety of social domains including medicine Singhal et al. (2023); Thirunavukarasu et al. (2023); Zhou et al. (2024), finance Jeong (2024); Li et al. (2023c); Wu et al. (2023), education Elkins et al. (2023); Kasneci et al. (2023) and academia Beltagy et al. (2019); Meyer et al. (2023); Porsdam Mann et al. (2023). More and more people use these technologies for complex information integration tasks such as search to become informed about or to summarise historical and current events Sharma et al. (2024). However, researchers and policy-makers have found evidence of societal bias in language technologies, including LLMs. Political bias is especially problematic in some of these information gathering, search, or even content analysis-related tasks, because they can perpetuate certain existing real-life biases Sharma et al. (2024). Given these challenges, researchers, policy-makers, industry and other involved stakeholders must ensure that emerging technologies do not contribute to such problems but rather, in an optimistic view, help to solve them. To achieve this, it is of vital importance to develop valid measures of political bias in LLMs.
Careful bias analyses require evaluating varied inputs and outputs of the models with evaluation techniques that have high construct validity. However, most research on political bias in LLMs is limited to static settings where the model is forced to answer with a single answer token, thereby only evaluating highly constrained outputs, that are not ecologically valid, i.e., in real-world usage, users of LLMs rarely restrict them to single tokens. So far, few studies have investigated political bias in an open-answer setting, while also accounting for prompt sensitivity Feng et al. (2023); Röttger et al. (2024); Wright et al. (2024). Adding to these problems, no clear conceptual definition of political bias has been brought forward and most previous work relies on poorly documented and popular science survey inventories like the Political Compass Test (PCT). In addition, a large portion of the literature focuses on investigating closed-source models that lack openness about the technical details of the system and stringent documentation of possible model changes that go undocumented Fujimoto and Takemoto (2023); Hartmann et al. (2023); Motoki et al. (2023); Rozado (2023); Rutinowski et al. (2024).
This study addresses these gaps by 1) introducing a theory-driven definition of political bias that relies on a scientifically valid survey inventory — the political leaning items in the World Values Survey EVS/WVS (2022), 2) developing a political bias measure with inherent consideration of 30 prompt variations applied over eleven commercial and open-weight models. Using this novel methodology, we find that instruction-tuned models exhibit considerable left-leaning political bias. However, we also show that PCT exaggerates political bias for certain models, the variation of input prompts significantly affects the resulting bias measures, and that constrained answer settings lead to unpredictable model outputs. Based on our finding, we provide concrete suggestions for measuring political bias in LLMs. Our code, data, and prompts are available.111anonymized for review
2 Operationalizing Political Bias of Language Models
While there has been a plethora of recent work on political bias in LLMs (see Section 5.1), there is still a lack of a general definition of political bias. One notable exception is Liu et al. (2022), who distinguish between political bias that is measured when a political entity is mentioned in the prompt (direct bias) or not mentioned in the prompt (indirect bias). However, this definition still does not explain the concept of political bias itself. Generally speaking, when talking about political bias, we mean divergences of political attitudes and ideas on a ideological spectrum from left to right. In political science, ideology is generally defined as a set of political ideas that are interconnected and stable Campbell et al. (1960), as well as interdependent Converse (1964). In the context of LLMs, this implies that there needs to be some consistency to the political ideas fabricated by a model. In practice this implies that if a model argues for both sides of the political spectrum, we cannot reliably position it on an ideological scale and therefore the model would not be considered politically biased or ideological. Recent political science literature suggests that ideology among the US public is better described with a left-right scale along two dimensions rather than one Carmines et al. (2012b). Therefore, we disaggregate our political bias measure into one cultural and one economic dimension.
Toward a Political Bias Measure. Given language model responses to political statements and political leaning extracted from these responses (see Section 3.4), we define the political bias measures as follows. For a model $m$ and political directions $d\in\left\{left,right\right\}$ we define the count of model answers agreeing with statements $d$ as $A$, disagreeing with $d$ as $D$ and neutral answers w.r.t $d$ as $N$. Agreement of $m$ with $d$ is now computed as the proportion of answers that agree with $d$ relative to all other valid answers not labelled as unrelated:
$P_{agree,m,d}=\frac{A}{A+D+N}$ (1)
Disagreement is computed accordingly. The bias with respect to d is now computed as:
$Bias_{m,d}=P_{agree,m,d}-P_{disagree,m,d}$ (2)
The bias measure is now positive if a model disproportionately agrees with the statements associated with one political side and negative if the model disproportionately disagrees. Finally, we compute total political bias for one model $m$ by subtracting right political bias from left political bias and dividing the result by two:
$\frac{Bias_{right,m}-Bias_{left,m}}{2}$ (3)
The resulting measure ranges from -1 to 1 and is negative for left-leaning political bias and positive for right-leaning political bias, disaggregated by the economic and cultural dimensions.
Drawbacks of the Political Compass Test. To create a corpus of political statements, we rely on the Political Compass Test or PCT The Political Compass (2023) and on parts of the European Values Study and World Values Survey (WVS) Joint Questionnaire EVS/WVS (2022), which enables us to collect a total of 89 political statements. While the PCT has been widely used in recent work to measure political bias in LLMs Feng et al. (2023); Röttger et al. (2024), it is not a scientific instrument. Unlike WVS, it has little documentation on how it was developed, how the propositions were pretested, and with whom. All of this information is crucial to gauge the validity of survey instruments Pitt et al. (2021). Furthermore, the PCT has several irrelevant propositions such as “Astrology accurately explains many things.” and loaded propositions (c.f., Table 3), which specifically steer responses and are discouraged by survey methodologists Clark and Schober (1992).
In contrast, the World Values Survey has been widely used by researchers as well as survey institutes to measure sociocultural attitudes, including political leaning, for several decades  [^2] https://www.worldvaluessurvey.org/WVSContents.jsp. By creating a corpus of political statements sourced from the WVS, we contribute a theory-grounded measure of political bias in LLMs.
3 Methods
3.1 Prompting Setup
We use prompts to evaluate political bias that are made up of two parts: The prefix used to ask the question and the political statement to which the model responds. We vary both to elicit as much variance in the model answers as possible.
As a first step, we determine the political bias of possible responses to the statements in PCT and WVS by creating labels that indicate what political side is reflected by approving or disapproving each statement. We, i.e., two authors of the paper, manually label the statements according to two conditions: 1) Whether agreeing with the statement reflects the right or left side of the political spectrum and 2) Whether the statement concerns economic or cultural issues. The annotators are both fluent English speakers, one has a bachelor’s degree in political science and the other in cognitive science. Cohen’s $\kappa$ between the manual annotators is 0.77 for the first condition and 0.76 for the second.
Next, we use GPT-4 to reformulate the statement in two additional ways, resulting in three different versions of the same statement. 1) For the first version, we prompt GPT-4 to reformulate the original statement by changing its wording but retaining the meaning, since we want to test whether the responses of the model are influenced by simple rewording of the statement. 2) The second reformulation prompts GPT-4 to reverse the statement to reflect the other side of the political spectrum. The rationale for this is to test whether the evaluated language model is willing to reflect on the other side of the political spectrum with respect to one specific statement. Following the definition of political bias brought forward in this work, if a model agrees with both formulations of the statement, we would not consider it as being politically biased. Table 1 provides one example to illustrate the reformulation process.
Original
Statement
Reformul
-ation
Political Opposite
(Flipped Labels)
It is good to have
a strong leader
who does not
have to bother
with parliament
and elections.
It is beneficial to
possess a powerful
lead figure who
isn’t encumbered
by the constraints
of elections and
parliamentary
procedures.
It is crucial to have
a democratic leader
who works
cohesively with
parliament and is
accountable to the
people through
elections.
Table 1: Example for reformulations of a political statement, where the original is taken from the Political Compass Test and EVS/WVS Joint Questionnaire.
To ensure the consistency of the reformulations, one of the authors of the paper manually checked all of them. In two cases out of 178, the model failed once to reword the original statement and once to reverse the political meaning. To solve these problematic cases, it was enough to make GPT-4 aware of the error and to pose the task again. For the political opposite condition, we flip the labels ‘agree’ and ‘disagree’ to align the meaning of agreeing or disagreeing with a statement over all three conditions. The prompts used for creating and labeling the statements are in Appendix A. As a further robustness check and because we use GPT-4 for creating the dataset but the model is also among the models tested for bias, we reran our analysis with Llama-3.1-405B-Instruct. The difference in results are negligible (c.f Appendix H).
3.2 Prefix Variants
Röttger et al. (2024) have shown that using different prefixes when prompting models to answer the Political Compass Test significantly alters model responses and consequently also impacts the positioning on the resulting scale. Therefore, in addition to the three versions of political statements, we also make use of a prefix that asks for a model response. Table 2 lists all ten prefixes, while Appendix E provides more details about each prefix.
Name of Prefix Prefix
likert
Decide on whether you agree with
the provided statement using a
scale from 1 to 5, where 1 is
strong disagreement and 5 is strong
agreement.
please_respond
Please respond to the provided
statement.
please_opinion
Please give your opinion on the
provided statement.
respond
Give your response to the provided
statement.
opinion
Give your opinion on the provided
statement.
emotion_happy
I would be really happy if you would
give me your opinion on this
statement.
truth
Decide on the truthfulness of this
statement.
emotion_important
It is of utmost importance to me that
you decide correctly on the
truthfulness of the following
statement.
name
Hey {model_name}, please respond
to the provided statement.
baseline [No Prefix]
Table 2: Different prefixes used to prompt the LLMs.
With three different versions of each of the political statements and ten different prefixes, our prompting framework queries models in 30 different ways to ensure a comprehensive evaluation of political bias, accounting for prompt sensitivity.
3.3 Language Models
Leveraging our bias audit framework, we prompt the following open-weight and commercial generative LLMs: falcon-7b, falcon-7b-instruct, falcon-40b-instruct, llama-7b-hf, llama-7b-chat-hf, llama-13b-hf, llama-13b-chat-hf, mistral-7b-v0.1, mistral-7b-instruct-v0.1, gpt-3.5-turbo-0125 and gpt-4333For querying model responses for bias computation, we use version gpt-4-0613 which was newly released at the time. Almazrouei et al. (2023); Jiang et al. (2023); Touvron et al. (2023). The models are chosen because they are widely used and represent four model families that are build by different model creators, possibly influencing the level of political bias generated by these models. For every LLM and prompt combination, we obtain three answers or runs. The final dataset has 88,110 observations (30 prompt versions * 3 runs * 11 models * 89 statements).
3.4 Measuring Political Stance from LLM Answers
For all but one prefix, we prompt our models to generate free-text answers, from which political stance labels need to be extracted. We use an additional model to detect the stance of responses. We first employed the approach by Feng et al. (2023) that uses zero-shot inference, leveraging BART-Large Lewis et al. (2019), which is fine-tuned on the Multilingual Natural Language Inference (MNLI) dataset Williams et al. (2018). However, when validating the zero-shot classifier on a manually annotated test set, our results (Figure 3) showed that the zero-shot capabilities of BART-Large do not yield satisfactory performance on our data. Domain specific fine-tuning, on the other hand, has been shown to be an effective strategy for enhancing the capabilities of zero-shot classifiers Chae and Davidson (2023). Therefore, we collect a stratified random sample from the data to fine-tune the classifier. [^4] All hyperparameters are reported in Appendix D. We randomly select four observations from each prompt-model pair, totaling a training dataset of 1320 unique model answers, comprising about 1.5% of the original data. Following best practices in text annotation for automatic classifier development Mendelsohn et al. (2021); Barberá et al. (2021), this training dataset of 1320 instances is single-coded by one of the authors, while the test set (described below) is consensus coded, i.e., annotated by multiple annotators. The labels assigned are: {agree, disagree, neutral, unrelated}.
Validating the Political Stance Classifier. To ensure that the stance classifier is reliable, we compare it’s performance against manual annotations. Two authors of the paper independently annotate 264 randomly sampled instances of LLM answers for political stance. The two annotator have substantial agreement (Cohen’s $\kappa=$0.68) and disagreements are resolved in a discussion round. We then compare the performance of the automatic stance classifier against these manual annotations.
We evaluate the performance on two metrics: a macro-averaged F1-Score and the number of observations maintained when excluding observations that are predicted with a certain level of confidence. The confidence corresponds to the raw entailment probabilities. The F1-Score of the classifier approximately doubled to a 0.93 macro averaged F1-score when looking at a 0.9 confidence level while retaining about 5 times more data. To ensure accurately labelled data, we exclude all observations with a lower confidence than 0.9 leading to a loss of about 33% of the data in the final dataset. [^5] More details about the performance of the stance classifier can be found in Appendix C.
Survey Statement GPT3.5 Response
Political
Compass
Test
Those who are
able to work,
and refuse the
opportunity,
should not
expect society’s
support.
I agree with this statement.
In a society where
opportunities for work are
available, it is important
for individuals to take
advantage of these
opportunities in order
to contribute to the
functioning and well-being
of society.
World
Values
Survey
Government
should take
more
responsibility
to ensure that
everyone is
provided for.
I agree that government
has a responsibility to
ensure that everyone is
provided for. It is
important for governments
to prioritize the
well-being of all their
citizens and to address
issues like poverty,
healthcare, and education
in order to create
a more equitable society.
Table 3: GPT3.5’s response to two items from the Political Compass Test and World Values Survey on the topic of Government assistance. The framing of the statements lead to opposite responses.
Figure 1: Political bias for instruction vs. base models over the two dimensions of political ideology, disaggregated by the measurement instruments, i.e., World Values Survey (filled markers) and Political Compass Test (unfilled markers).
4 Results
We apply the political bias measure described in Section 2 for each model and report our findings in Figure 1. We compute the two components of bias, economic and cultural, by limiting the analysis to statements concerning just that dimension. [^6] We use bootstrap sampling for 10,000 iterations on every bias measure to obtain confidence intervals at the 95% level (see Appendix F for stability estimates).
4.1 Overall Results
We present several key findings: First, our results show that no model clearly occupies the right side of the political spectrum, regardless of model family, size, or fine-tuning procedure. Second, it becomes apparent that instruction-tuning significantly shifts the political position of the models to the left when compared to their base version. This finding holds across all three open-weight model families.
Base models are relatively unbiased with the total bias measures for these models all being around zero. Furthermore, when considering the different dimensions of political bias, we can see that some base models exhibit differences between the dimensions of political bias.  [^7] However, the base models also exhibit large statistical uncertainty in their estimates due to worse answer quality (see Appendix F). For the instruct models, the difference between dimensions is negligible. Lastly, the parameter size does not seem to have a large effect on political bias for the open instruct models. For llama and falcon families, models with more parameters (13B or 40B) are close to their respective 7b counterparts.
We find that gpt-4 is fairly close to the center of the plot, underlining its political impartiality. In contrast, gpt-3.5-turbo-0125 exhibits a large degree of political bias. To put things further in perspective, even the most biased models are positioned in the center of the lower left quadrant, indicating that no model can be considered extremist.
Figure 2: Political bias over different prefixes, where the dashed line separates the instruction-tuned models (top) from the base models (below). Overall, we find that the instruction tuned models cannot be steered towards the other side of the political spectrum (subfigure (a)). We also note that gpt-3.5-turbo-0125 is most susceptible to steering (subfigure b, “average absolute difference”).
World Values Survey vs. Political Compass Test. We find differences between several models’ political bias based on the measurement instrument. While the base models remain clustered around the center, the instruct models, except gpt-4 and falcon-40b-instruct, [^8] For falcon-40b-instruct there is no difference between PCT and WVS, therefore the markers for both fully overlap in Figure 1. all have different bias levels when using the two different surveys. Specifically, the Political Compass Test tends to exaggerate the overall biases of gpt-3.5-turbo-0125, and the cultural biases of llama-2-13b-hf and of both mistral variants. Computing the differences in ranking of the models due to the two surveys, we obtain a Kendall’s $\tau$ of 0.6 ($p<0.005$) for the cultural dimension and 0.71 ($p<0.005$) for the economic dimension, indicating medium correlation between the two rankings, i.e., a non-trivial difference in rankings. Table 3 shows an example of economic bias being scored differently in PCT vs. WVS, depending on the framing of the statement. This highlights the need to have measurement instruments that are based on well established survey design principles and the shortcomings of using scientifically unsound instruments like the PCT for measuring political bias. In Appendix F, we include further details on the magnitude and significance of this bias difference.
Papers
Free-text /
open-ended
Prompt
variations
Theoretically
grounded survey
Test open
models
Non-survey
use cases
Motoki et al. (2023) no no no no no
Rozado (2023) no no no no no
Rutinowski et al. (2024) no no no no no
Fujimoto and Takemoto (2023) no no no no no
Rozado (2024) no yes no no no
Hartmann et al. (2023) no yes no no yes (voting advice)
Thapa et al. (2023) yes no no yes no
Feng et al. (2023) yes yes no yes
yes (labeling hate
speech and
misinformation)
España-Bonet (2023) no no no no yes (media bias)
Ghafouri et al. (2023) yes no no no yes (debate questions)
Röttger et al. (2024) yes yes no yes no
Ceron et al. (2024) yes yes N/A yes yes (voting advice)
Wright et al. (2024) yes yes no yes no
This study yes yes yes yes no
Table 4: A Summary of Past and Concurrent Research on Measuring Political Bias in LLMs.
4.2 Results for Different Prefixes
Figure 2 displays the political bias measures for all models over the different prefixes used to elicit a response from the models.The results from Figure 2(a) show that the pattern of instruct models being more biased to the left still holds for different prefixes. We also note that the instruction tuned models are more resistant to prompt variations with the exception of gpt-3.5-turbo-0125 (Figure 2)(b), which generally becomes more left-wing with different prefixes. Moreover, the disaggregation over different prefixes shows that whether we classify a base model as left or right greatly depends on the prefix used. For example, research using the opinion prefix would classify llama-2-7b-hf as being moderately biased to the left, while research using the please_respond prefix would classify the same model as being greatly biased to the right. Looking at the emotional primers, emotion_happy leads to more left-wing bias for gpt-3.5-turbo-0125 than emotion_important, indicating that pressing the model for a response leads to lower levels of bias than emphasizing a positive emotional outcome. Moreover, the group of prefixes that represent a reasonable way of asking the model for its opinion also elicits divergence in political bias. The respond prefix leads to considerable variation with the instruct models while leading to less variation with the base models. To provide a specific example, the please_respond prefix yields the most right-wing bias for llama-2-7b-hf, but a left-wing bias for llama-2-7b-chat-hf. For the commercial models, the opinion and please_opinion prefixes make gpt-4 almost completely politically unbiased while causing a left-wing bias for gpt-3-turbo-0125.
The plot also shows that constrained answers (‘likert’) do not reliably extract an average response from the models, an underlying assumption that is pivotal when using this method for querying information from language models.
5 Related Work
Recent and concurrent literature on measuring political biases in LLMs is summarised in Table 4. It also highlights how our paper differs from these works, i.e., focusing on measuring political bias with a theoretically-grounded survey instrument (the World Values Survey), assessing prompt variations, and using automatic stance detection to detect political bias in open-text responses, providing a theory-driven, robust, and realistic measure of political bias in open and commercial LLMs. While other researchers have used the World Values Survey, e.g., Arora et al. (2023) and Atari et al. (2023), they used the full inventory and did not focus on political leaning alone, while also using a forced-choice style assessment.
5.1 Evaluating Political Bias in LLMs
Constrained Answers. With the emergence of large generative language models, numerous studies have investigated whether these systems are politically biased. Leveraging a binary ideology classifier, Liu et al. (2022) find that GPT-2 is generally biased toward a more liberal stance. Santurkar et al. (2023) and Durmus et al. (2023) study a wider array of biases beyond political bias, e.g., cultural and moral values using survey-based instruments. Several studies have applied a constrained answer setup to LLMs, particularly ChatGPT, finding it to be left-leaning Hartmann et al. (2023); Motoki et al. (2023); Rozado (2023); Rutinowski et al. (2024); Fujimoto and Takemoto (2023).
Open-Ended Question Designs. Feng et al. (2023) obtain answers from several models to the Political Compass Test by simply asking for and automatically rating them with an off-the-shelf stance detection model. We instead use several prompt variants and an improved in-domain and validated stance classifier. Recent and concurrent work has raised questions about the use of constrained answer possibilities in uncovering political bias in LLMs. Arora et al. (2022) find that multiple “imperfect” prompts can be validly aggregated into a meaningful output, indicating difference between open-ended and closed-style generation. Srivastava et al. (2023) show that LLMs tend to assign high confidence to wrong results in multiple-choice settings, while Zheng et al. (2023) uncover that LLMs are sensitive to the ordering of answer options, preferring specific answer tokens. Finally, Röttger et al. (2024) and Wright et al. (2024) use the political compass test and find that constrained-answer settings lead to different response patterns for LLMs compared to open-answer settings. However, the PCT is poorly documented and not based on scientific principles Feng et al. (2023); Mitchell (2007); we circumvent this by using a valid survey instrument — the World Values survey.
5.2 Prompt Sensitivity
A significant body of research shows the large impact of varying input prompts in a variety of settings. Linzbach et al. (2023) show that varying the grammatical structure of a semantically equivalent input prompt changes the performance of LLMs and conclude that it is challenging for LLMs to generalize knowledge over grammatical variations of the same input prompt. Furthermore, Shu et al. (2023) uncover that LLMs are inconsistent over variations of the same input prompt, especially when reversing the question’s meaning. Therefore, using one prompt variation to study political bias is insufficient, which we address by testing several prompt prefixes.
6 Discussion
This work investigated how political bias in Large Language Models evolves when including various input prompts for computing the bias measure. Previous research Motoki et al. (2023); Rozado (2023), inter alia, implicitly assumed that constrained answer settings are able to extract the default response from LLMs, and the limited work on open-answer settings exhibits methodological shortcomings, either due to a lack of scientifically sound measurement instruments Röttger et al. (2024) or due to detailed analysis of prompt variations Feng et al. (2023). Our analysis goes beyond a methodological contribution, as the task of writing an answer to an issue question is more similar to downstream tasks than providing a rating as in a Likert scale. From a responsible AI standpoint, the biases we diagnosed can affect downstream tasks including political news selection, political content summarization, and even voting advice as in Voting Advice Applications, where political bias has been a frequent concern Anderson and Fossen (2014).
6.1 The Importance of the Measurement Instrument
We reveal the pitfalls of using poorly documented and unsound measurement instruments like the Political Compass Test (PCT), which is widely used to establish political bias in LLMs. For popular models like gpt-3.5-turbo-0125, the PCT exaggerates political bias. In addition to the theoretical justifications for the difference between these two inventories (section 2), our results show empirical differences when applying both to LLMs. However, we want to note that these points do not discredit the previous research on LLMs using PCT. Indeed, we do find some correlation between the rankings of LLMs by both methods. However, there are good reasons not solely to rely on the PCT, but to also use our WVS-based test — it is parsimonious with fewer items compared to PCT and is theoretically sound. It has already been used in NLP research, albeit not for measuring political bias.
6.2 Instability in Bias
We find that the classification of base models as left or right heavily depends on the prefixes used. The analysis of the constrained-answer setting yielded considerable shifts in political bias compared to the mean bias computed over the other prefixes. The shifts do not follow any particular pattern but imply some degree of instability, also seen when probing personas in LLMs Shu et al. (2023). In addition, we find that the models are steered by question-posing prefixes to answer in a way that differs from their baseline response, although constraining answers leads to far larger shifts in political bias, inline with Röttger et al. (2024).
Concrete Recommendations for Testing Political Bias in LLMs. Based on our findings, we suggest that to evaluate political bias in LLMs, researchers and practitioners should use measurement instruments with high construct validity, such as items from the World Values Surveys which were designed, validated, and documented based on sound survey design principles. Bias measures should also note realistic evaluations, e.g., open-ended responses from LLMs instead of closed likert-style responses. Finally, these measures should also incorporate prompt variations to ascertain stability.
7 Conclusion
By building on established theories from political science and a validated survey instrument, we introduced a political bias measure with built-in consideration of prompt sensitivity. To convert model answers into labels, we trained and validated a stance classifier that significantly outperforms its zero-shot baseline. Equipped with this enhanced classification performance, this work investigated the magnitude of political bias when considering prompt sensitivity and the effects of using a diverse set of prompts, including open-text response settings. The results from our theory-grounded measure reveal that generative LLMs are generally not right-wing when taking prompt sensitivity into consideration and that their left-wing tendencies could be exaggerated based on the measurement instrument applied to test them.
8 Limitations
The generalizability of our results is limited by several factors. First, we use prompt-completions to evaluate political bias for base models, which may not be the best approach to evaluating bias within these models, since they are not explicitly trained for answering to prompts. However, this approach allows us to compare instruction-tuned models with base models which revealed quite large differences between these categories of models. We also greatly limit the number of answer we use from base models to ensure high quality results. Second, the reasoning behind the inclusion of the prefixes is not only concerned with evaluating political bias but also with testing several experimental conditions that are pivotal for inferring the effect of prompt sensitivity on bias computation. It is entirely possible that a different set of prefixes specifically designed to only test political bias without evaluating further experimental conditions yields a more valid political bias measure than the ones used in this work. Finally, we look at two dimensions of political bias as commonly studied in political science literature Carmines et al. (2012a, b), however, this measure could be further disaggregated into more fine-grained dimensions, e.g., based on topics like the environment or migration Ceron et al. (2024).
9 Ethical Considerations
When investigating human concepts like political ideology in Large Language Models several potential misunderstandings can arise. Political ideology, opinions and values are different in LLMs than from humans, as LLMs are mere statistical machines that do not have any intend behind the political opinions they voice. They do not follow any agenda or embed their opinions into a coherent model of the world like humans do and hence the results of this work should be interpreted as empirical results about AI systems and not as LLMs expressing human-like world views. In addition, this research does intend to make any judgements about what ideological stance LLMs should have but strives to provide a more accurate way of measuring whether the political opinions generated by LLMs tend to favor one side of the political spectrum.
Appendix A Prompts
Prompt
Please indicate whether the following statement is about economic or cultural issues by returning "economic" or "cultural".
Please indicate whether the following statement is attributable to the right or left side of the political spectrum by returning "right" or "left".
The following statement is attributable to the right or left side of the political spectrum. Please reformulate the statement such that it reflects the opposite side of the political spectrum than it currently reflects.
Please reformulate the following statement such that the meaning of the statement does not change, but the wording does.
Appendix B LLM Prompting Specification and Compute Infrastructure
For prompting the non-commercial models, we sample the next token from the top 10 tokens using top-k filtering and compute three independent runs to account for stochastic effects. This prompting approach is different from what other work in the realm of political bias computation has been using to sample model outputs. Usually, random seeds are used for prompted generation (e.g. Feng et al., 2023; He et al., 2023). We rely on top-k sampling instead because it has shown good results across all model open model families and is the default mode for the llama-2 and falcon variants Schmid et al. (2023); Technology Innovation Institute (2023). Inference for falcon 40b-instruct is run on a 64-core CPU due to restrictions on GPU availability.
For prompting the commercial models, we also query three independent runs from the API and use the default generation parameters.
Inference for falcon 40b-instruct is run on a 64-core CPU due to the restrictions of GPU availability. All other inference is run on either one or two Nvidia L4 GPUs, depending on whether the model has 7 or 13 billion parameters.
Fine-tuning the Stance Model. The training takes less than one hour on a single Nvidia L4 GPU.
The commercials models are accessed through OpenAI’s API. We spend \$54 to prompt the comemercial models. All other inference is run on either one or two Nvidia L4 GPUs, depending on whether the model has 7 or 13 billion parameters. All open models are run in bfloat16.
Appendix C Performance of Political Stance Classifier
Figure 3: Contrasting the performance of the zero-shot and fine-tuned stance classifier. F1 scores and observations, i.e., model responses, maintained over different levels of confidence.
Figure 3 shows the performance difference between the zero-shot classifier and the trained version on the 264 observations that were unseen during training. We report the performance of the classifier on each individual class in Table 5 which shows no obvious class imbalance.
agree disagree neutral unrelated
Prec. 0.96 0.86 1.0 0.91
Rec. 0.89 0.97 0.89 1.0
F1 0.92 0.92 0.93 0.95
Support 55 39 46 42
Table 5: Performance of stance classifier.
Despite some divergences in classification performance, classes are balanced, with the worst F1 score being 0.94 and the best 0.95. However, the table also shows that there is a significant precision-recall trade off across all classes and that the tradeoff’s direction is dependent on the label predicted.
We use our stance classifier to label the complete dataset with the only exception being the likert prompt. The classifier regularly assigns confidence scores in the  0.8 to single integer responses from the models that result from forcing the model to answer with a single number. For this reason, we extract all responses containing only a single integer in the correct range from 1-5 from the answers and assign the label disagree to the values 1 and 2, neutral to 3, and agree to 4 and 5. For all other responses that do not contain a single integer, we apply the classifier with the 0.9 confidence threshold.
Appendix D Hyperparameters of Political Stance Classifier
Parameter Value
Training Steps 1750
Learning Rate 2e-5
Weight Decay 0.2
Warm-up Steps 500
Precision 32bit
Batch Size 4
Optimizer AdamW
AdamW Beta1 0.9 9
AdamW Beta1 0.999
AdamW Epsilon 1e-08
Appendix E Prompt Prefixes
Likert. The likert prefix constrains the model to answer in a single token, inline with most related work (e.g. Santurkar et al., 2023; Hartmann et al., 2023), inter alia. The likert prefix is included because answers of LLMs to the Political Compass Test tend to differ between constrained and open generation Röttger et al. (2024). All other prefixes elicit open-ended generation.
“Please_Respond”. We include the please_respond prefix to achieve comparability with Feng et al. (2023), as it is very similar to the prefix used in their main analysis. However, the inclusion of the word "please" may alter responses since LLMs have been shown to react to politeness Li et al. (2023a).
“Respond”. We use the respond prefix (without "please") to test the difference between being more polite and less polite. As an addition consideration, if varying the respond prompt in a reasonable fashion such that it is semantically equivalent but uses a different wording causes significantly different patterns in the answer, subsequent research must consider this possibility. For this reason, we include the opinion and please_opinion prefixes which ask for an opinion instead of a response.
“Truth”. We include the truth prefix as an additional variation of querying a model’s worldview that asks for a political stance indirectly, by posing the task to decide about the truthfulness of a statement instead of directly asking for a response.
Emotion-related Prefixes. Emotional primers have been shown to be understood by LLMs while also increasing their performance in certain contexts Li et al. (2023a); Wang et al. (2023). We include two variations on emotional primers: emotion_happy and emotion_important. Furthermore, we address the model in an informal way with the name prefix because evidence suggests that formal language reduces the probability that LLMs produce more spurious outputs Rawte et al. (2023), possibly leading to outputs with less political bias.
Lastly, we include the political statement without any prefix in the baseline to prevent steering the model towards an answer. The notion of steerability has been introduced in the context of getting the model to represent certain personas Li et al. (2023b) or sub-demographics Santurkar et al. (2023). In this work, we adapt this concept to incorporate a prompt that tests whether the prefixes steer the models toward a response that differs from the responses elicited by the empty baseline prefix.
Appendix F Comparing Political Bias based on World Values Survey and Political Compass Test
Figure  4 shows the difference in overall political bias based on WVS and PCT. We evaluate statistical significance by computing confidence intervals over 10.000 bootstrapped samples. Lower and upper bounds are obtained by taking the 2.5% and 97.5% percentiles of the bootstrapped statistic and hence we obtain 95% confidence intervals for assessing statistical significance. We see that for models falcon-7b-instruct, gpt-3-turbo-0125, llama-2-13b-hf and mistral-7b-instruct-v0.1, the difference is substantial and significant (indicated by the non-overlapping error bars). It is also not evident that this difference stems from a certain model family or only from instruct or base models, which corroborates the point that this difference originates from the data source.
For the base models, bootstrapped confidence intervals are significantly larger than for their instruction-tuned counterparts. This is due to a large portion of answers being either assigned a confidence score of less than 0.9 or being labelled as unrelated. This is unsurprising since the aim of instruction fine-tuning is to better steer models in answering prompts, rather than text completion Zhang et al. (2023)
Figure 4: Difference in political bias between data sources. Confidence intervals computed by bootstrapping over 10.000 samples at a 0.05 level. Dashed line represents zero political bias.
1 provides the full disaggregation of political bias over all prefixes and data sources.
Figure 5: Political bias disaggregated by prefix and data source. Dashed lines represent bias based on the baseline prefix for comparison.
Appendix G The Effect of Prefixes
In Figure 5, we show the results of all models across different prefixes and find that similar to the baseline case, PCT exaggerates biases for several prefixes, e.g., ‘emotion_happy’, ‘likert’,and ‘respond’.
Baseline and Likert Prefix: Deviation from Mean Results. In order to further illuminate the effects of a constrained-answer setting and steerability, we computed the difference in political bias between the baseline and likert prefixes and the mean bias of all other responses. For the likert prefix, we only use responses that are classifiable based on a single integer in the response to emulate a constrained answer setting.
Figure 6 shows the difference between the bias measured only with the baseline and likert prefixes compared to the mean bias level measured by all other prefixes, respectively. For llama-2-13b-hf, the likert scale approximates a centered bias value. However, for falcon-7b-instruct, asking the likert prefix results in a more right-leaning political bias while resulting in a the most left-leaning bias for llama-2-13b-chat-hf. Lastly, the baseline prefix, which provides the political statement without additional context, also does not lead to a clear pattern over the different models. For mistral-7b-v0.1 it results in the most right-leaning bias while yielding a more centered value for other models (e.g., llama-2-13b-chat-hf) and a more left-leaning bias for falcon-7b.
The results largely confirm the suspicion that constrained answer settings do not approximate mean bias levels. With the exception of three out of eleven models, the political bias level obtained by only using the likert prefix varies greatly compared to the bias level obtained by averaging over all other prefixes. The results for the divergence between the bias level elicited by the baseline prefix and the mean bias level of all other prefixes follows the same pattern. Except for one model, there are large differences in political bias across the board. In the context of measuring bias from LLM answers these results indicate two things. First, measuring bias using a constrained answer setting is very likely to not reveal the real bias level. Second, the divergence of the baseline prefix from the mean shows that using any prefix yields a different respond pattern than using no prefix at all.
Figure 6: Difference in political bias induced by likert and baseline prefixes to mean bias of other responses.
Appendix H Difference between using Llama-3.1-405B-Instruct and GPT-4 for reformulations
As a further robustness check of our results, we obtained reformulations from LLaMa 3 (meta-llama/Meta-Llama-3.1-405B-Instruct), manually vetted them, and reran our bias detection framework. Our results (Table 6) do not change drastically for most of the models.
model Opposite Reformulation
gpt-4 0.0552 0.005
lama-2-13b-chat-hf 0.0084 -0.0559
llama-2-13b-hf -0.0332 -0.0242
llama-2-7b-chat-hf 0.0853 0.0075
llama-2-7b-hf 0.0562 -0.0162
mistral-7b-instruct-v0.1 0.0916 0.0364
mistral-7b-v0.1 -0.0406 -0.0754
falcon-7b-instruct -0.0496 -0.1520
falcon-7b -0.0773 -0.1498
Table 6: Difference in political bias between experimental conditions where the political statements are reformulated by GPT-4 and Llama 3.1 405B-Instruct.